OpenStudy (anonymous):

Solve. 1/36 = 6^(x-3)

OpenStudy (lgbasallote):

6^-2 = 6^(x-3) -2 = x - 3 x = 1

OpenStudy (experimentx):

6^-2=6^(x-3) -2=x-3 x=1

OpenStudy (anonymous):

Try putting the non-powered things in the same terms (or base)